# ClimaOoD: Improving Anomaly Segmentation via Physically Realistic Synthetic Data

### v1

@@ 1 Introduction
Anomaly segmentation aims to detect out-of-distribution (OoD) objects not predefined, serving as a complementary capability to semantic segmentation methods[liu2023residual, nayal2023rba, lis2019detecting, rai2024mask2anomaly, blum2021fishyscapes]. This capability is crucial for safety-critical applications such as autonomous driving, where models must respond reliably to unexpected hazards like fallen cargo, construction equipment, or animals on the road[djeumou2025onemodel, chen2025solve, xu2025vlmad, DiBiase2021pixelwise]. Despite notable advances in deep segmentation networks, their performance in open-world environments remains highly constrained by the availability and diversity of anomaly data[chan2021segmentmeifyoucan, xia2025opensetanomalysegmentationcomplex, blum2021fishyscapes, pinggera2016lost, chan2021entropy, delić2024outlierdetectionensemblinguncertainty, lis2019detecting].
In the real world, anomalous events are inherently rare and unpredictable, which makes collecting large-scale, high-quality anomaly data for training anomaly segmentation models extremely difficult[zhang2023improving, xia2025opensetanomalysegmentationcomplex, nekrasov2025spotting, blum2021fishyscapes]. Existing benchmarks often focus on urban streets under clear weather, overlooking complex environmental conditions (e.g., tunnels, highways, rain, fog, or snow)[pinggera2016lost, blum2021fishyscapes, chan2021segmentmeifyoucan, lis2019detecting] (see Figure 1.a). Consequently, models trained on these datasets struggle to generalize when faced with unseen weather or novel environments[rai2024mask2anomaly, zhang2023improving].
To alleviate the scarcity of anomaly data, recent state-of-the-art methods have explored two main paradigms of synthetic data generation. The first is the copy-paste approach [liu2023residual, zhao2024segment, nayal2023rba], which crops objects from external datasets (e.g., COCO [lin2014microsoft], ADE20K [zhou2019semantic]) and pastes them into driving scenes (e.g., Cityscapes [cordts2016cityscapes]) to simulate OoD events. The second leverages text-to-image diffusion models [loiseau2024reliability, xia2025opensetanomalysegmentationcomplex], where masked regions in driving images are filled with anomalous objects guided by textual prompts. While both approaches enrich anomaly diversity, they remain limited in complementary ways: copy-paste methods struggles to maintain contextual consistency, often causing significant discrepancies in brightness and color between pasted objects and backgrounds. Meanwhile, text-to-image diffusion-based approaches lack physical realism, leading to implausible object placement regarding positional coherence and scale (see Figure 1.b). These issues create domain gaps between synthetic training data and real-world application scenarios, ultimately compromising model robustness and generalization capability.
@@
Built upon ClimaDrive, we construct ClimaOoD, a large-scale benchmark collected to evaluate anomaly segmentation robustness in open-world conditions. ClimaOoD synthesizes over 10,000 image-mask pairs across multiple weather conditions (clear, rain, fog, snow, cloudy, nighttime) for training. After careful curation, we create a refined test set of 1,200 image-mask pairs, covering six distinct driving scenarios (e.g., city streets, highways, tunnels) and 93 anomaly types. In comparison to existing anomaly segmentation benchmarks such as LostAndFound [pinggera2016lost] (1 landform, 9 anomaly types), Fishyscapes [blum2021fishyscapes] (1 landform, 7 anomalies), and SMIYC-RoadAnomaly21 [chan2021segmentmeifyoucan] (4 landforms, 26 anomalies), ClimaOoD provides substantially broader coverage with 6 landforms, 93 anomaly categories, and 6 distinct weather conditions. This wide diversity ensures more robust training for anomaly segmentation models under varied open-world conditions.
@@ 2.1 Anomaly Segmentation Benchmarks in Autonomous Driving
Anomaly segmentation relies heavily on high-quality and diverse datasets[zhang2023improving, zhao2024segment, lis2019detecting, shoeb2025outofdistributionsegmentationautonomousdriving]. Yet real-world anomaly collection remains difficult—anomalous events are rare, unpredictable, and highly variable across environments[Liu2025OoDDINOAM, Zheng2025SegmentingOA]. This scarcity limits model generalization to diverse out-of-distribution (OoD) scenarios[kim2023sanflow, sun2025anomalyany, zhou2023pad].
Early benchmarks such as LostAndFound[pinggera2016lost] and Fishyscapes[blum2021fishyscapes] focus mainly on clear-weather urban scenes, offering limited anomaly and scene diversity. RoadAnomaly[lis2019detecting] and SMIYC[chan2021segmentmeifyoucan] expand scene types and weather conditions but remain small in scale. ComsAmy[xia2025opensetanomalysegmentationcomplex] increases scenario and weather coverage yet contains only 468 images, restricting robust evaluation. Overall, existing datasets are too limited or too small to reflect the complexity of real-world driving environments.
@@ 2.2 Synthetic Data Generation Methods
In anomaly segmentation, generating synthetic data has become a crucial strategy to overcome the limitations of real-world datasets[Tian2022pixelwise, Lis2020detecting, Vojir2021road, shoeb2025outofdistributionsegmentationautonomousdriving]. Two primary methods have been widely adopted: the copy-paste approach and text-to-image diffusion models.
The copy-paste method inserts objects from external datasets into driving scenes[Tian2022pixelwise, liu2023residual, nayal2023rba, Liu2025OoDDINOAM], as used in Fishyscapes-Static[blum2021fishyscapes] and StreetHazards[steinhardt2019benchmark]. While it enables easy anomaly diversification, it often breaks contextual consistency—pasted objects frequently mismatch scene geometry, scale, or lighting. Text-to-image diffusion synthesizes anomalies directly via prompts[nie2024compositional, ohanyan2024zero, liang2024rich, loiseau2024reliability]. Loiseau et al.[loiseau2024reliability], for instance, generate anomalies from randomly sampled boxes. Despite improved visual diversity, these methods lack physical grounding, leading to spatially or geometrically misaligned anomalies.
@@
LostAndFound [pinggera2016lost] 0.12% 1203 1 9 clear
Fishyscapes LAF [blum2021fishyscapes] 0.23% 373 1 7 clear
SMIYC-RoadObstacle21 [chan2021segmentmeifyoucan] 0.12% 327 2 31 clear, frosty, night
RoadAnomaly [lis2019detecting] 9.85% 60 3 21 clear
SMIYC-RoadAnomaly21 [chan2021segmentmeifyoucan] 13.83% 100 4 26 clear
@@ 3.1 Generation Framework: ClimaDrive
Scene attributes $\alpha=\{\alpha^{\text{weather}},\alpha^{\text{scene}},\alpha^{\text{time}}\}$ are derived from BDD100K[yu2020bdd100k] metadata and fused with a caption $c$ from BLIP [li2022blip] to form $p=f(c,\alpha)$. Stable Diffusion [rombach2022highresolution] with ControlNet [lv2023controlnet] is employed, where ControlNet is fine-tuned using $S_{\text{sem}}$ as structural guidance and $p$ as the linguistic condition. This preserves geometric alignment while achieving realistic variation across weather and lighting conditions.
@@
Additionally, grounding-based anomaly masks $M_{\text{ano}}$ are generated using the Grounding-SAM method[zhu2024grounded], and these masks are further refined using a lightweight noise-denoise process to improve boundary smoothness and coherence, ensuring that the generated anomalies align well with the scene structure.
@@ 4 Experiment
Mask2Anomaly[rai2024mask2anomaly] FishyScape LAF 95.41 69.46 9.31 96.28 74.15 7.86
@@
Methods. These four methods represent two main paradigms in out-of-distribution(OoD) anomaly segmentation. Outlier-exposure methods (RPL[liu2023residual], Mask2Anomaly[rai2024mask2anomaly]) aim to directly highlight abnormal regions: RPL[liu2023residual] learns residual patterns to increase pixel-level sensitivity, while Mask2Anomaly[rai2024mask2anomaly] formulates anomaly detection as mask classification with masked attention. Uncertainty-based methods (RbA[nayal2023rba], UNO[delić2024outlierdetectionensemblinguncertainty]) calibrate anomaly scores by modeling prediction confidence: RbA[nayal2023rba] suppresses overconfident in-distribution logits, and UNO[delić2024outlierdetectionensemblinguncertainty] adopts a K+1K{+}1 classifier with an explicit outlier class to refine anomaly scoring.
@@ 4.1 Training OoD Segmentation Models
To assess the impact of training with the ClimaOoD dataset, we compare the performance of four state-of-the-art OoD segmentation models when trained with ClimaOoD versus original datasets. The models are evaluated on four benchmark datasets (Fishyscapes LAF[blum2021fishyscapes], RoadAnomaly[lis2019detecting], RoadAnomaly21[chan2021segmentmeifyoucan], and RoadObstacle21[chan2021segmentmeifyoucan]) to examine the generalization capability under diverse driving scenarios and weather conditions.
Results and Analysis. Table 2 summarizes the performance improvements when training with ClimaOoD. Across all methods, training with ClimaOoD results in notable gains in AUROC, AP, and FPR95. For example, RbA[nayal2023rba] achieves an increase in AUROC from 97.19 to 97.85 on Fishyscapes LAF, and AP improves from 78.27 to 81.32 on RoadAnomaly. Similarly, Mask2Anomaly[rai2024mask2anomaly] sees improvements in both AUROC (from 95.41 to 96.28) and AP (from 69.46 to 74.15) across the datasets.
In particular, training with ClimaOoD enhances the robustness of these models, with FPR95 decreasing in most cases (e.g., RbA[nayal2023rba] drops from 3.97 to 3.52 on Fishyscapes LAF). These results demonstrate that ClimaOoD provides valuable, diverse training data that leads to better generalization and performance under a wide range of OoD conditions. The improvements highlight the significance of training with realistic, weather-diverse data for improving OoD segmentation in open-world environments.
@@ 4.2 Benchmarking OoD Segmentation Models
Scenario Weather RPL [liu2023residual] RbA[nayal2023rba] Mask2Anomaly[rai2024mask2anomaly] UNO[delić2024outlierdetectionensemblinguncertainty]
@@
Although UNO[delić2024outlierdetectionensemblinguncertainty] and RPL[liu2023residual] achieve the strongest overall results, all methods show clear robustness gaps in complex environments. Unlike prior benchmarks such as Fishyscapes[blum2021fishyscapes]—restricted to clear-weather urban scenes—ClimaOoD exposes failure cases across diverse weather and lighting, providing a more realistic and comprehensive evaluation of OoD segmentation performance.
@@ 4.4 Technical Analysis
Table 5 presents an analysis of the key components in the ClimaDrive framework. The goal of this analysis is to understand the contribution of each component to the overall performance in generating realistic OoD data. The performance is evaluated using four key metrics: Fréchet Inception Distance (FID), Learned Perceptual Image Patch Similarity (LPIPS), and two Pearson Correlation Coefficients (PCC) based on entropy and max-logit scores, which measure the alignment between generated data and real-world OoD data from the SMIYC-RoadAnomaly21[chan2021segmentmeifyoucan] benchmark, due to the similarity in object scale to our synthetic data. Specifically, entropy and max-logit are used to assess pixel-level uncertainty in model predictions.
@@
Additional, we compare SD1.5[rombach2022high], SD2[stabilityai_sd2_inpainting], and SDXL[SDXL:2023] as inpainting backbones under the same training loss. Table 6 shows SDXL gives the best quality (FID=164.8, LPIPS=0.263) but is computationally heavy. SD2 achieves nearly the same realism with 40% fewer GFLOPs and is adopted as default.

### v2

@@ 1 Introduction
Anomaly segmentation aims to detect out-of-distribution (OoD) objects not predefined, serving as a complementary capability to semantic segmentation methods[18, 22, 16, 27, 3]. This capability is crucial for safety-critical applications such as autonomous driving, where models must respond reliably to unexpected hazards like fallen cargo, construction equipment, or animals on the road[11, 7, 36, 10]. Despite notable advances in deep segmentation networks, their performance in open-world environments remains highly constrained by the availability and diversity of anomaly data[6, 35, 3, 26, 5, 9, 16].
In the real world, anomalous events are inherently rare and unpredictable, which makes collecting large-scale, high-quality anomaly data for training anomaly segmentation models extremely difficult[39, 35, 23, 3]. Existing benchmarks often focus on urban streets under clear weather, overlooking complex environmental conditions (e.g., tunnels, highways, rain, fog, or snow)[26, 3, 6, 16] (see Figure 1.a). Consequently, models trained on these datasets struggle to generalize when faced with unseen weather or novel environments[27, 39].
To alleviate the scarcity of anomaly data, recent state-of-the-art methods have explored two main paradigms of synthetic data generation. The first is the copy-paste approach [18, 40, 22], which crops objects from external datasets (e.g., COCO [15], ADE20K [42]) and pastes them into driving scenes (e.g., Cityscapes [8]) to simulate OoD events. The second leverages text-to-image diffusion models [20, 35], where masked regions in driving images are filled with anomalous objects guided by textual prompts. While both approaches enrich anomaly diversity, they remain limited in complementary ways: copy-paste methods struggles to maintain contextual consistency, often causing significant discrepancies in brightness and color between pasted objects and backgrounds. Meanwhile, text-to-image diffusion-based approaches lack physical realism, leading to implausible object placement regarding positional coherence and scale (see Figure 1.b). These issues create domain gaps between synthetic training data and real-world application scenarios, ultimately compromising model robustness and generalization capability.
@@
Built upon ClimaDrive, we construct ClimaOoD, a large-scale benchmark collected to evaluate anomaly segmentation robustness in open-world conditions. ClimaOoD synthesizes over 10,000 image-mask pairs across multiple weather conditions (clear, rain, fog, snow, cloudy, nighttime) for training. After careful curation, we create a refined test set of 1,200 image-mask pairs, covering six distinct driving scenarios (e.g., city streets, highways, tunnels) and 93 anomaly types. In comparison to existing anomaly segmentation benchmarks such as LostAndFound [26] (1 landform, 9 anomaly types), Fishyscapes [3] (1 landform, 7 anomalies), and SMIYC-RoadAnomaly21 [6] (4 landforms, 26 anomalies), ClimaOoD provides substantially broader coverage with 6 landforms, 93 anomaly categories, and 6 distinct weather conditions. This wide diversity ensures more robust training for anomaly segmentation models under varied open-world conditions.
@@ 2.1 Anomaly Segmentation Benchmarks in Autonomous Driving
Anomaly segmentation relies heavily on high-quality and diverse datasets[39, 40, 16, 30]. Yet real-world anomaly collection remains difficult—anomalous events are rare, unpredictable, and highly variable across environments[19, 41]. This scarcity limits model generalization to diverse out-of-distribution (OoD) scenarios[12, 32, 43].
Early benchmarks such as LostAndFound[26] and Fishyscapes[3] focus mainly on clear-weather urban scenes, offering limited anomaly and scene diversity. RoadAnomaly[16] and SMIYC[6] expand scene types and weather conditions but remain small in scale. ComsAmy[35] increases scenario and weather coverage yet contains only 468 images, restricting robust evaluation. Overall, existing datasets are too limited or too small to reflect the complexity of real-world driving environments.
@@ 2.2 Synthetic Data Generation Methods
In anomaly segmentation, generating synthetic data has become a crucial strategy to overcome the limitations of real-world datasets[33, 17, 34, 30]. Two primary methods have been widely adopted: the copy-paste approach and text-to-image diffusion models.
The copy-paste method inserts objects from external datasets into driving scenes[33, 18, 22, 19], as used in Fishyscapes-Static[3] and StreetHazards[31]. While it enables easy anomaly diversification, it often breaks contextual consistency—pasted objects frequently mismatch scene geometry, scale, or lighting. Text-to-image diffusion synthesizes anomalies directly via prompts[24, 25, 14, 20]. Loiseau et al.[20], for instance, generate anomalies from randomly sampled boxes. Despite improved visual diversity, these methods lack physical grounding, leading to spatially or geometrically misaligned anomalies.
@@
LostAndFound [26] 0.12% 1203 1 9 clear
Fishyscapes LAF [3] 0.23% 373 1 7 clear
SMIYC-RoadObstacle21 [6] 0.12% 327 2 31 clear, frosty, night
RoadAnomaly [16] 9.85% 60 3 21 clear
SMIYC-RoadAnomaly21 [6] 13.83% 100 4 26 clear
@@ 3.1 Generation Framework: ClimaDrive
Scene attributes $\alpha=\{\alpha^{\text{weather}},\alpha^{\text{scene}},\alpha^{\text{time}}\}$ are derived from BDD100K[37] metadata and fused with a caption $c$ from BLIP [13] to form $p=f(c,\alpha)$. Stable Diffusion [29] with ControlNet [21] is employed, where ControlNet is fine-tuned using $S_{\text{sem}}$ as structural guidance and $p$ as the linguistic condition. This preserves geometric alignment while achieving realistic variation across weather and lighting conditions.
@@
Additionally, grounding-based anomaly masks $M_{\text{ano}}$ are generated using the Grounding-SAM method[44], and these masks are further refined using a lightweight noise-denoise process to improve boundary smoothness and coherence, ensuring that the generated anomalies align well with the scene structure.
@@ 4 Experiment
Mask2Anomaly[27] FishyScape LAF 95.41 69.46 9.31 96.28 74.15 7.86
@@
Methods. These four methods represent two main paradigms in out-of-distribution(OoD) anomaly segmentation. Outlier-exposure methods (RPL[18], Mask2Anomaly[27]) aim to directly highlight abnormal regions: RPL[18] learns residual patterns to increase pixel-level sensitivity, while Mask2Anomaly[27] formulates anomaly detection as mask classification with masked attention. Uncertainty-based methods (RbA[22], UNO[9]) calibrate anomaly scores by modeling prediction confidence: RbA[22] suppresses overconfident in-distribution logits, and UNO[9] adopts a K+1K{+}1 classifier with an explicit outlier class to refine anomaly scoring.
@@ 4.1 Training OoD Segmentation Models
To assess the impact of training with the ClimaOoD dataset, we compare the performance of four state-of-the-art OoD segmentation models when trained with ClimaOoD versus original datasets. The models are evaluated on four benchmark datasets (Fishyscapes LAF[3], RoadAnomaly[16], RoadAnomaly21[6], and RoadObstacle21[6]) to examine the generalization capability under diverse driving scenarios and weather conditions.
Results and Analysis. Table 2 summarizes the performance improvements when training with ClimaOoD. Across all methods, training with ClimaOoD results in notable gains in AUROC, AP, and FPR95. For example, RbA[22] achieves an increase in AUROC from 97.19 to 97.85 on Fishyscapes LAF, and AP improves from 78.27 to 81.32 on RoadAnomaly. Similarly, Mask2Anomaly[27] sees improvements in both AUROC (from 95.41 to 96.28) and AP (from 69.46 to 74.15) across the datasets.
In particular, training with ClimaOoD enhances the robustness of these models, with FPR95 decreasing in most cases (e.g., RbA[22] drops from 3.97 to 3.52 on Fishyscapes LAF). These results demonstrate that ClimaOoD provides valuable, diverse training data that leads to better generalization and performance under a wide range of OoD conditions. The improvements highlight the significance of training with realistic, weather-diverse data for improving OoD segmentation in open-world environments.
@@ 4.2 Benchmarking OoD Segmentation Models
Scenario Weather RPL [18] RbA[22] Mask2Anomaly[27] UNO[9]
@@
Although UNO[9] and RPL[18] achieve the strongest overall results, all methods show clear robustness gaps in complex environments. Unlike prior benchmarks such as Fishyscapes[3]—restricted to clear-weather urban scenes—ClimaOoD exposes failure cases across diverse weather and lighting, providing a more realistic and comprehensive evaluation of OoD segmentation performance.
@@ 4.4 Technical Analysis
Table 5 presents an analysis of the key components in the ClimaDrive framework. The goal of this analysis is to understand the contribution of each component to the overall performance in generating realistic OoD data. The performance is evaluated using four key metrics: Fréchet Inception Distance (FID), Learned Perceptual Image Patch Similarity (LPIPS), and two Pearson Correlation Coefficients (PCC) based on entropy and max-logit scores, which measure the alignment between generated data and real-world OoD data from the SMIYC-RoadAnomaly21[6] benchmark, due to the similarity in object scale to our synthetic data. Specifically, entropy and max-logit are used to assess pixel-level uncertainty in model predictions.
@@
Additional, we compare SD1.5[28], SD2[2], and SDXL[1] as inpainting backbones under the same training loss. Table 6 shows SDXL gives the best quality (FID=164.8, LPIPS=0.263) but is computationally heavy. SD2 achieves nearly the same realism with 40% fewer GFLOPs and is adopted as default.